Gordon Corporation produced 10,000 digital watches in the current year. Variable costs are $8 per watch. Overhead assigned is $2.25 per watch. A supplier offers the watches for $9.50 each. Gordon's production manager reports the incremental overhead is $1.25 per watch. Gordon should:

  • Continue making the watches as an additional $1.50 per watch would be incurred if bought from the supplier.

  • Buy the watches as they would save $0.75 per watch.

  • Buy the watches as they would save $1.50 per watch.

  • Buy the watches as they would save $1.75 per watch.

  • Continue making the watches as an additional $0.25 per watch would be incurred if bought from the supplier.
Answer #1
Total relevant costs of making watches = 8+1.25 = $9.25
Additional cost of buying watches = 9.50-9.25 = $0.25
Gordon should Continue making the watches as an additional $0.25 per watch would be incurred if bought from the supplier.
Option E is correct